### v3

# sample.bib

Abstract
In the past two decades, extensive research has been conducted on the (co)homology of various models of random simplicial complexes. So far, it has always been examined merely as a list of groups. This paper expands upon this by describing both the ring structure and the Steenrod-algebra structure of the cohomology of the multiparametric models. For the lower model, we prove that the ring structure is always a.a.s trivial, while, for certain parameters, the Steenrod-algebra a.a.s acts non-trivially. This reveals that complex multi-dimensional topological structures appear as subcomplexes of this model. In contrast, we improve upon a result of Farber and Nowik, and assert that the cohomology of the upper multiparametric model is a.a.s concentrated in a single dimension.
1 Introduction
1.1 Background
Random simplicial complexes and their topological properties have been an active area of study in recent years (see, for instance, [HandbookDnCG3, chapter 22], or [Bobrowski_2022]), where many results are of the form ”property $Q$ is satisfied asymptotically almost surely” or a.a.s, that is with probability approaching 1 as the number of vertexes tends to $\infty$. Two models that have received extensive study are the clique complex of a random graph from $G(n,p)$ ([KAHLE09]), and the Linial-Meshulam-Wallach model, first appearing in [MeshulamWallach] (the 2-dimensional version first appearing in [LinialMeshulam]), comprised of the complete $m$-skeleton on $n$ vertexes, and a random set of $m+1$-simplices.
Definition 1.1.
Take a random hypergraph $X(n;p_{1},p_{2},...)$ composed of the (independently) chosen facets of each dimension, where facets of dimension $i$ are chosen with probability $p_{i}$. We may then form two distributions on simplicial complexes: $\underline{X}(n;p_{1},p_{2},...)$ and $\overline{X}(n;p_{1},p_{2},...)$.
$\underline{X}(n;p_{1},p_{2},...)$ will be the maximal simplicial complex contained in $X(n;p_{1},p_{2},...)$ and is called the Lower Multiparametric Model or the Lower Model for short.
$\overline{X}(n;p_{1},p_{2},...)$ will be called the Upper Multiparametric Model (or the Upper Model for short), and is taken to be the minimal simplicial complex containing the hypergraph. These were first introduced in [CostaFarber] and [FarberMeadNowik19].
These may be thought of in the following way: $\underline{X}(n;p_{1},p_{2},...)$ is the distribution on simplicial complexes with $n$ vertexes, where each edge is included independently with probability $p_{1}$, each triangle whose boundary formed in the first step is filled independently with probability $p_{2}$ and so on. $\overline{X}(n;p_{1},p_{2},...)$, may be thought of in the same way, but without requiring the boundary to exist before attaching a simplex. We will also write $X(n;\alpha_{1},\alpha_{2},...)$ to mean that $p_{k}=n^{-\alpha_{k}}$ (and similarly $\overline{X},\underline{X}$).
$\underline{X}(n;p_{1},p_{2},...)$ interpolates between the clique complex and the Linial-Meshulam-Wallach model (with parameters $(p,1,1,...),(1,1,...,1,p_{r},0,...)$ respectively), while $\overline{X}(n;p_{1},p_{2},...)$ is its obvious ”dual”.
In both the clique and Linial-Meshulam-Wallach models cohomology is concentrated in one dimension. However, in [Fowler19], Fowler studied the asymptotic behavior of the cohomology of the lower multiparametric model with coefficients in $\mathbb{Q}$, and found that multipule cohomologies can be non-zero at the same time:
Theorem 1.2 (Fowler).
Denote:
$S_{1}^{k}(\{\alpha_{i}\}_{i})=\Sigma_{i=1}^{k+1}\binom{k+1}{i}\alpha_{i}\qquad S_{2}^{k}(\{\alpha_{i}\}_{i})=\Sigma_{i=1}^{k}\binom{k+2}{i+1}\alpha_{i}$
Then:
  • If $S_{1}^{k}(\{\alpha_{i}\}_{i})<1$, then $H^{k}(\underline{X},\mathbb{Q})=0$ a.a.s (see Definition 2.9).
  • If $S_{2}^{k}(\{\alpha_{i}\}_{i})>k+2$, then $H^{k}(\underline{X},\mathbb{Z})=0$ a.a.s.
  • If $S_{1}^{k}(\{\alpha_{i}\}_{i})\geq 1$ and $S_{2}^{k}(\{\alpha_{i}\}_{i})<k+2$ and all $\alpha_{i}$ are positive, then $H^{k}(\underline{X},\mathbb{Q})\neq 0$ a.a.s.
The situation in $\overline{X}$ was more subtle. We quote from [FarberNowik22], using a notation convention that for a fixed complex $A$ we denote by $E(A\subset X)$ the expected amount of its appearances as a subcomplex of a random complex $X$, or simply $E(A)$ if $X$ is clear from context.:
Definition 1.3.
In $\overline{X}(n;\alpha_{1},\alpha_{2},...)$:
  • $\beta_{i}:=i+1-\alpha_{i}\approx log_{n}E(\Delta^{i}\subset X)$, $\beta:=max\{\beta_{i}\}$, $l=\lfloor\beta\rfloor$
  • $\gamma_{k}:=max_{i\geq k}\{\beta_{i}\}\approx log_{n}E(\Delta^{k}\subset Y)$
  • $\nu_{k}:=2\gamma_{k}-k\geq log_{n}(b_{k})$ (k’th betti number), $l^{\prime}:=max\{k|\nu_{k}\geq 0\}\leq\lfloor 2\beta\rfloor\leq 2l+1$
  • $e_{k}:=\gamma_{k}-k=\nu_{k}-\gamma_{k}<0$ where the inequality is when $k>l$. $e_{k+1}\leq e_{k}-1$
(some facts included for intuition and convenience).
Formulated in this notation, Farber and Nowik proved the following:
Theorem 1.4.
For $\alpha$’s such that $\beta\notin\mathbb{Z}$, $\overline{X}(n;\alpha_{1},\alpha_{2},...)$ a.a.s has a full $(l-1)$-skeleton, collapses onto dimension $l^{\prime}$, and has $dim(H_{l}(\overline{X};\mathbb{Q}))=\theta(n^{\beta})$. Further, a.a.s $dim(H_{k}(\overline{X};\mathbb{Q}))\leq\omega(n)n^{\nu_{k}}$ for $k>l$ and any $\omega(n)\xrightarrow{n\to\infty}\infty$. So $H_{k}(\overline{X}),l<k<l^{\prime}$ was suspected, but not known, to be non-trivial.
These two articles exhibit a phenomenon in the existing literature- most if not all papers in the field discuss cohomology merely as a list of Betty numbers. Cohomology (with ring coefficients), on the other hand, is a graded ring itself, with the multiplication operation being the cup product. Moreover, cohomology with coefficients in $\mathbb{F}_{p}$ also has a structure of a steenrod algebra ([AT] chapter 4.L) which is the algebra of stable cohomology operations (that is natural transformations between cohomology functors that commute with the equivalence $H^{*}(\Sigma X)=H^{*-1}(X)$) over those fields, generated by the steenrod squares $Sq^{i}:H^{*}\to H^{*+i}$ in the case of $p=2$, or the steenrod powers $P^{i}:H^{*}\to H^{*+2i(p-1)}$ and the Bockstein homomorphism $\beta:H^{*}\to H^{*+1}$ in the case of an odd $p$. These additional structures are usfull in distinguishing between spaces with otherwise identical cohomology. The following are classic examples:
Example 1.5.
Let $X=S^{2}\vee S^{1}\vee S^{1}$ and $Y=\mathbb{T}^{2}$, the 2-dimensional torus. Both have identical cohomology groups $H^{0}=\mathbb{Z},H^{1}=\mathbb{Z}^{2},H^{2}=\mathbb{Z}$. However, $H^{*}(X)$ has a trivial multiplicative structure, whereas in $H^{*}(Y)$ the product of the two generators of $H^{1}$ equals the generator of $H^{2}$, proving that $X$ and $Y$ are not homotopy equivalent.
Example 1.6.
Let us compare $\mathbb{C}P^{2}$ and $Y=S^{2}\vee S^{4}$. Again these spaces have identical comohology over any ring (due to the cell structure), yet the former has a non-trivial cup product while the later does not.
Furthermore, in $\mathbb{C}P^{2}$ the generator of $H^{2}$ squares to the generator of $H^{4}$, which in $\mathbb{F}_{2}$ coefficients coincides with $Sq^{2}$, which unlike the cup product is stable. Thus, the spaces $\Sigma\mathbb{C}P^{2}$ and $\Sigma Y=S^{3}\vee S^{5}$, which no longer can be distinguished by the cup product structure (as it is trivial for both), may still be distinguished by the non-triviality of $Sq^{2}$ on the former.
[…]
The goal of this paper is to determine whether this structure is trivial or not for the multiparametric models. To the authors knowledge, this is the first time that non-trivial multidimensional phenomena are studied in stochastic topology.
1.2 Results
Definition 1.7.
For a topological space $X$, the cup length of $X$ is the maximal amount of elements $\{a_{i}\}|a_{i}\in H^{n_{i}}(X),\forall i\ n_{i}>0$ such that $\Pi_{i}a_{i}\neq 0$.
Theorem 1.8.
Given a list of positive real numbers $\{\alpha_{i}\}_{i=1}^{\infty}$ that satisfy $S_{1}^{k}(\{\alpha_{i}\}_{i})\neq 1$ for all $k\in\mathbb{N}$, the family $\underline{X}(n;\{\alpha_{i}\}_{i})$ a.a.s has cup length of 1 over $\mathbb{Q}$. Moreover, the image of the cup product in $H^{2l}$ is trivial over any ring of coefficients, where $l\in\mathbb{N}$ is any number s.t $S^{l}_{1}(\{\alpha_{i}\}_{i})>1$.
Thus, for a full measure of the $\alpha$’s, no non-trivial cup product exists. This is proven by noticing that the requirement of having both $H^{k},H^{2k}$ be non-zero is very restrictive in this model. These restrictions entail limited possible homotopy types of strongly connected components (see Definition 2.2), in which the cup and steenrod operations are concentrated (in a sense made concrete in Lemma 2.7).
Unlike the cup product, steenrod operations are stable and thus each one requires only a range of cohomologies of a fixed length to exist (as opposed to the upper dimension being twice the lower), and thus the restriction is less severe.
Theorem 1.9.
Fix a non-zero element $\sigma$ of the steenrod algebra. Then there exist $D\in\mathbb{N}$ and a set $\{\alpha_{i}\}_{i=1}^{D}\subset\mathbb{R}_{\geq 0}^{D}$ with non empty interior such that $\sigma$ is a.a.s non-zero as a map on $H^{*}(\underline{X}(n;\{\alpha_{i}\}),\mathbb{Z}/p)$.
To prove this, we construct a variation of a simplicial model of the suspension operation (see 5.2). This variation “increases the odds” that a complex to which it is applied appears as a subcomplex. If we apply this variation sufficiently many times, the complex appears as a subcomplex of $\underline{X}(n;\{\alpha_{i}\})$ a.a.s. This construction preserves several desirable properties of the complex, including having a non-zero Steenrod operation and strong connectivity. Thus, for some range of $\alpha$’s, $\sigma$ is non-zero on the complex.
This indicates that the lower model might have highly non-trivial topological structure. The same can not be said for the upper model- simplexes of dimension greater than $l$ can not form sufficiently complex structures with each other, and as a result:
Theorem 1.10.
For $\{\alpha_{i}\}_{i}$ such that $\beta\not\in\mathbb{Z}$, $\overline{X}(n;\alpha_{1},...)$ a.a.s collapses onto dimension $l$.
1.3 Structure of the paper
In Section 2, we gather definitions and results useful in the analysis of $\overline{X},\underline{X}$ and the topology of simplicial complexes in general. The proof of Theorem 1.8 is presented in Section 3 and Section 4. Section 5 is devoted to the proof of Theorem 1.9, and Section 6 is devoted to the proof of Theorem 1.10. Lastly, Section 7 details problems left open by the present manuscript.
2 Preliminaries
2.1 Topology & Probability
Definition 2.1.
A simplicial complex is called pure $d$-dimensional if any simplex in the complex is contained in a $d$ dimensional simplex. Note that this implyes that $d$ is the maximal dimension.
Definition 2.2.
For a simplicial complex define a relation $\sim$ on its $d$-dimensional simplexes by having $\sigma\sim\tau$ if they share a $d-1$-dimensional face. This is obviously reflexive and symmetric. Complete this relation under transitivity and call it $\sim^{\prime}$. A Strong Connectivity Component is an equivalence class of $d$-dimensional faces under this relation. A pure-$d$-dimensional simplicial complex is called Strongly Connected if all of its $d$-dimensional faces are in the same component. A simplicial complex where all maximal simplexes are of dimension greater or equal to $d$ is called Strongly Connected with respect to dimension $d$ if its $d$-skeleton is strongly connected.
Definition 2.3.
Let $a,b$ be two cochains of Dimension $k,l$ respectively. Then the cup product $a\cup b$ is defined on a simplex $\Delta^{k+l}=(v_{0},...,v_{k+l+1})$ by
$(a\cup b)(v_{0},...,v_{k+l+1})=a(v_{0},...,v_{k+1})\cdot b(v_{k+1},...,v_{k+l+1})$
This produces a graded product structure on cohomology (see [AT]). The important thing for us is that the value of the cup product on a particular simplex depends only on the values of the multiplied cochains on the faces of that simplex.
Remark 2.4.
In this paper we consider a cup product non-trivial only when there exist two cocycles of positive dimension whose product is non-zero in cohomology, or equivilently when the cup-length is strictly greater than 1. For a connected space $X$, $H^{0}(X;R)=R$, and multiplication by a 0-dimensional cocycle is equivalent to multiplication by a scalar. Thus omitting the condition of the cocycles being of positive dimension would make the question uninteresting. The disconnected case is similar.
Definition 2.5.
Steenrod operations are defined on a simplicial level by performing the cup-i product of a cochain with itself. A definition of the cup-i product first appears in [Cup_i_Steenrod] for cohomology with $\mathbb{Z}_{2}$ coefficients. Later on in [Steenrod53-1] and [Steenrod53-2] Steenrod defined similar operations for $\mathbb{Z}/p$ for odd $p$. The definitions are quite involved and will not be explicitly stated in this paper. Partially for that reason most topologists have transitioned to using more conceptual definitions stemming from abstract considerations (the one for $\mathbb{Z}/2$ coefficients may be seen in the last chapter of [AT]).
The important thing for us is that the value of $Sq^{i}(c)$ (for $\mathbb{Z}/2$ coefficients) or $P^{i}(c)$ and $\beta(c)$ (for $\mathbb{Z}/p$ coefficients) on a simplex $s$ depends only on the values of the co-cycle $c$ on faces of $s$. A modern treatment of the subject (including a method for constructing explicit formulas for the $\mathbb{Z}/p$ case) may be found in [Kaufmann_2021].
For a steenrod operation $\sigma$, we denote by $deg(\sigma)$ the amount by which it increases dimension.
Lemma 2.6.
Let $i:A\to B$ be an inclusion of simplicial complexes. Then:
  1. 1.
    If $c$ is a cocycle/coboundery on $B$, then so is $i^{*}c$.
  2. 2.
    If $c^{\prime}$ is a coboundery on $A$, then $\emptyset\neq(i^{*})^{-1}(c^{\prime})\subset Im(\delta_{B})$. In other words, $c^{\prime}$ can be extended to a coboundary on $B$.
The proof of Lemma 2.6 is simple and is left to the reader. We now turn to the key insight that enables the results in this paper:
Lemma 2.7.
  1. 1.
    For a simplicial complex $X$, $H^{d}(X)$, $Im(\vee:H^{*}(X)\times H^{*}(X)\to H^{d}(X))$ and steenrod operations landing in $H^{d}(X)$ are non-zero only if there exists a $d$-dimensional strong-connectivity component on which they are non-zero. If $X$ has $d$ as the maximal dimension, this is an if and only if for $H^{d}(X)$.
  2. 2.
[…]
Proof.
  1. 1.
    Denote the strong connectivity components of $X$ by $\{X_{i}\}_{i}$. Define $f:\bigsqcup_{i}X_{i}\to X$ to be the map which restricts to the inclusion on each $X_{i}$. Observe that by Lemma 2.6 $f^{*}$ does not create new coboundaries in dimension $d$, we conclude that $f^{*}$ is injective on $H^{d}$. If $X$ is $d$ dimensional then it is surjective as well since $f^{*}$ is an isomorphism on $d$-dimensional cochains and all of them are cocycles.
    Assume $0\neq a\vee b=c\in H^{d}(X)$. By the above there exists $X_{i}$ such that $c|_{X_{i}}\neq 0\in H^{d}(X_{i})$. $a|_{X_{i}},b|_{X_{i}}$ are cocycles, and since $\vee$ is natural we conclude $a|_{X_{i}}\vee b|_{X_{i}}=c|_{X_{i}}$, meaning the cup product on $X_{i}$ is non-trivial. The proof for steenrod operations is identical.
  2. 2.
    The claim about $r,s$ is obvious, and the second claim follows from naturality of the cup product and steenrod operations.
Corollary 2.8.
For any simplicial complex $X$, the non-triviality of any of $H^{d}(X)$, $\vee:H^{*}(X)\times H^{*}(X)\to H^{d}(X)$ or $\sigma:H^{d-deg(\sigma)}(X)\to H^{d}(X)$ implies its non-triviality for $sk^{d}(X)$. Further, to deduce the non-triviality of $H^{d}$, $\vee$ or $\sigma$ on the entire complex from that on a strong connectivity component, we need only check this cocycle condition. In particular, only the addition of simplexes of dimensions $dim(p)+1,dim(q)+1$ or $dim(r)+1$ (or $dim(s)+1$) might affect this.
Definition 2.9.
Let $D(n;\vec{p}(n))$ be a family of distributions depending on a natural number $n$ and perhaps other parameters. $D(n;\vec{p}(n))$ satisfies a property $q$ a.a.s (asymptotically almost surely) if $\underset{n\to\infty}{lim}P(q)=1$.
Example 2.10.
Let $G(n,p)$ be a distribution on graphs with $n$ vertexes, where each edge appears with probability $p$. A theorem by Erdős and Rényi states that $\frac{ln(n)}{n}$ is the threshold function for connectivity, meaning that for $p(n)=o(\frac{ln(n)}{n})$ the graph is a.a.s diconnected, and for $p(n)=\omega(\frac{ln(n)}{n})$ it is a.a.s connected.
2.2 General Results in the Lower Multiparametric Model
Proposition 2.11.
Let $X\in\underline{X}(n;\alpha_{1},...)$ be a random simplicial complex. Then a.a.s for any two $0\neq a,b\in H^{*}(X;\mathbb{Q})$ pure dimensional elements, $\frac{1}{2}\leq\frac{dim(a)}{dim(b)}\leq 2$.
Proof.
From Theorem 1.2 we know that the range where rational cohomology is non-zero is the $k$’s where
$\Sigma_{i=1}^{k+1}\binom{k+1}{i}\alpha_{i}\geq 1\land\Sigma_{i=1}^{k}\binom{k+2}{i+1}\alpha_{i}<k+2$
Take $k_{0}$ to be the minimal $k$ where the first condition is satisfied. Then inputting $k=2k_{0}+1$ in the second expression we get
$\Sigma_{i=1}^{2k_{0}+1}\binom{2k_{0}+3}{i+1}\alpha_{i}<2k_{0}+3\Longleftrightarrow\Sigma_{i=1}^{2k_{0}+1}\frac{\Pi_{j=1}^{i}(2k_{0}+3-j)}{(i+1)!}\alpha_{i}<1$
This contradicts the first expression, since it would mean in particular
$11\leq\Sigma_{i=1}^{k_{0}+1}\binom{k_{0}+1}{i}\alpha_{i}\leq\Sigma_{i=1}^{k_{0}+1}\frac{\Pi_{j=1}^{i}(2k_{0}+3-j)}{(i+1)!}\alpha_{i}<$ 1
where the first inequality follows from $\binom{k_{0}+1}{i}=\frac{\Pi_{j=1}^{i}(k_{0}-j+2)}{i!}<\frac{\Pi_{j=1}^{i}(2k_{0}-j+3)}{(i+1)!}$ which in turn is a consequence of $i+1<2^{i}<\Pi_{j=1}^{i}\frac{2k_{0}+3-j}{k_{0}+2-j}$.
This entails that in the lower-multiparametric model the ratio of dimensions in which the ($\mathbb{Q}$) cohomology is non-zero is at most 2. ∎
Corollary 2.12.
The only possible non-trivial cup product is of two cocycles in $H^{k}(X,\mathbb{Q})$ landing in $H^{2k}(X,\mathbb{Q})$.
Definition 2.13.
For a simplicial complex $X$ strongly connected w.r.t dimension $d$, an expansion operation is an operation which adds a simplex of dimension greater or equal to $d$, while keeping the component strongly connected. In order for this to happen, the added simplex will share a $d-1$ dimensional face with an existing face, in addition to possibly other faces.
Remark 2.14.
The non-pure dimensional case will only be relevant in Section 6.
In this paper, when discussing strongly connected components as resulting from expansion operations we refer to a budget as the expected amount of simplexes of dimension $d$ in the complex, and the cost of an operation as the factor by which the expected amount of the expanded complex is lower than the original. We will also often refer to the $\log_{n}$ of these by the same terms. Expansion operations fit into a ”growing process”, i.e series of subcomplexes $A_{1}\subset A_{2}\subset...A_{k}=A$, where $A_{1}$ is a full $d$-simplex, $A$ is any strongly connected $d$-complex, $A_{i+1}$ has exactly one $d$-simplex more than $A_{i}$ and all $\{A_{i}\}$ are strongly connected.
Example 2.15.
The following are the expansion operations for the pure 2-dimensional case:
Operation Cost Homotopical effect
$02\alpha_{1}+\alpha_{2}-1>$ 0 None
$12\alpha_{1}+\alpha_{2}>$ 1 Adds generator to $H^{1}$
$\alpha_{1}+\alpha_{2}>\frac{1}{2}$ None
$\alpha_{2}$ Reduces $H^{1}$ or increases $H^{2}$
where the costs are w.r.t the lower model and the inequalities follow from $S^{1}_{1}(\{\alpha_{i}\}_{i})>1$. The black and gray simplexes denote the complex before the expansion operation, while the red patterned simplexes denote new simplexes.
We now see how a tetrahedron is constructed:
$A$$B$$C$$A$$D$$B$$C$$A$$D$$B$$C$$A$$D$$B$$C$
So we begin from the triangle $ABC$ and a budget of $3-3\alpha_{1}-\alpha_{2}$, and proceed to add vertex $D$ via the first operation, then add triangle $ABD$ via the third operation (costing $\alpha_{1}+\alpha_{2}$), and finally ”close” the tetrahedron by adding triangle $BCD$ using the forth operation (costing $\alpha_{2}$). All in all, there remains of the budget $4-6\alpha_{1}-4\alpha_{2}$, which, assuming $S^{1}_{1}(\{\alpha_{i}\}_{i})\geq 1$ is less than or equal to $1-\alpha_{2}$, that is the expected number of subcomplexes isomorphic to the tetrahedron is $O(n^{1-\alpha_{2}})$.
Note that the vertexes $A,B,C,D$ are not particular vertexes of $X$, but are denoted in the isomorphism types for convenience only.
Remark 2.16.
It is at this point that we may give an interpretation to the expressions $S^{k}_{1}(\{\alpha_{i}\}_{i}),S^{k}_{2}(\{\alpha_{i}\}_{i})$ in the lower model. $S^{k}_{1}(\{\alpha_{i}\}_{i})\geq 1$ is the requirement that the expansion operation adding a vertex decreases the expected amount of appearances of the subcomplex, while $S^{k}_{2}(\{\alpha_{i}\}_{i})<k+2$ is the requirement that $k$-dimensional simplexes appear at all (asymptotically).
Note that where more than one choice of $d-1$ dimensional starting face is possible, we disregard the choice as unimportant. We generally think of a strongly connected complex as starting out of a single simplex and being added to by expansion operations one simplex at a time. There are multiple choices of both starting simplex and order of expansion operations performed. However, these generally do not matter for the following reasons:
Lemma 2.17.
For any strongly connected $d$-dimensional simplicial complex $X$, we may assume without loss of generality that all vertex-adding operations happened prior to any other operation.
Proof.
For the complex $X$ define a graph $G$ where a vertexes corresponds to the $d$-dimensional simplexes of $X$ and the edges correspond to the relation of sharing a $d-1$ dimensional face. $G$ being connected is equivalent to the definition of strong connectivity.
If a strongly connected complex with $v$ vertexes can be constructed only from vertex adding operations, then it will have $v-d$ $d$-dimensional faces. This is obviously an if and only if. If $X$ has more $d$-dimensional faces, we prove that it is always possible to remove one without removing vertices and while keeping the complex strongly connected.
Removing a face from $X$ is akin to removing a vertex from $G$ and all edges connected to that vertex. Assume for the sake of contradiction that there is no possible $d$-dimensional face to remove. Then for every vertex of $G$, removing it will result in a disconnected graph. But taking a spanning tree of $G$ and removing a leaf will always keep the graph connected, leading to a contradiction.
So there is always a $d$-dimensional face which it is possible to remove while keeping the complex strongly connected. If removing this face also removes a vertex, then this is the only face containing that vertex, meaning this face is a vertex addition operation. Thus removing this vertex and applying the induction hypothesis on the amount of vertices grants the result. ∎
Proposition 2.18.
Assume $S_{1}^{k}(\{\alpha_{i}\}_{i})\geq 1$. For a strongly connected $m$-dimensional complex $C$ denote by $C^{\prime}$ the result of some expansion operation adding $w$ edges and no vertexes to $C$, where $m>k$. Then $E(C)=\Omega(n^{\frac{w}{k+1}}E(C^{\prime}))$
Proof.
Any operation that adds an edge adds all simplecies supported on it. Therefore, the cost of any operation adding an edge will be higher than the operation adding only the edge and all simplecies supported on it. That cost is $\Sigma_{i=1}^{m-1}\binom{m-1}{i-1}\alpha_{i}$, as we only need to chose the vertexes distinct from the endpoints of the edge. We now compare coefficient-wise to our first inequality:
$\frac{\Pi_{j=1}^{i-1}(m-j)}{(i-1)!}=\binom{m-1}{i-1}\geq\frac{\binom{k+1}{i}}{k+1}=\frac{\Pi_{j=0}^{i-2}(k-j)}{i!}$
$\frac{\Pi_{j=1}^{i-1}(m-j)}{\Pi_{j=1}^{i-1}(k+1-j)}\geq 1\geq 1/i$
as $m\geq k+1$. Thus we see that the cost of adding an edge is at least $\frac{1}{k+1}$.
Moreover, assume we are preforming an operation adding $w$ edges simultaneously. This operation adds at least $\binom{m}{l}-\binom{m-w}{l}$ simplexes of dim $l$ (all missing edges share a vertex as the expansion operation is supported on a complete $m-1$ face. Thus it remains to chose $l$ other vertexes, one of which must be part of a missing edge). We wish to show this is bigger than $\frac{w}{m}\binom{m}{l}$.
$\binom{m}{l}-\binom{m-w}{l}\geq\frac{w}{m}\binom{m}{l}$
$\Pi_{j=0}^{l-1}(m-j)-\Pi_{j=0}^{l-1}(m-w-j)\geq w\Pi_{j=1}^{l-1}(m-j)$
$(m-w)\Pi_{j=1}^{l-1}(m-j)\geq\Pi_{j=0}^{l-1}(m-w-j)$
but this inequality is obvious as all multiplicands on the left are greater or equal than the ones on the right. $\frac{w}{m}\binom{m}{l}\geq\frac{w}{k+1}\binom{k+1}{l}$, finishing the proof. ∎
These expansion operations should thus be thought of as ”decreasing the likelihood” of a complex appearing as a subcomplex. This is actually true for all expansion operations:
Proposition 2.19.
Assume $S_{1}^{k}(\{\alpha_{i}\}_{i})\geq 1$ and $m\geq k$. For a strongly connected $m$-dimensional complex $D$ denote by $D^{\prime}$ the result of some expansion operation performed on $D$, adding (among others) an $l$-dimensional simplex. If either $S_{1}^{k}(\{\alpha_{i}\}_{i})>1$ or the operation doesn’t add vertexes and $\alpha_{l}>0$, then $E(D^{\prime})=o(E(D))$.
Proof.
$E(D)$ is, up to a constant factor, of the form $n^{f_{0}(D)-\Sigma_{i=1}^{\infty}f_{i}(D)\alpha_{i}}$, where $f_{i}(D)$ denotes the amount of $i$-dimensional faces of $D$. If the operation added a vertex, this costs at least $S^{m}_{1}(\{\alpha_{i}\}_{i})-1$, and since $S^{m}_{1}(\{\alpha_{i}\}_{i})\geq S^{k}_{1}(\{\alpha_{i}\}_{i})>1$ the result follows. Otherwise, we know that $f_{0}(D^{\prime})=f_{0}(D)$, $f_{l}(D^{\prime})\geq f_{l}(D)$ and that $f_{l}(D^{\prime})>f_{l}(D)$, and so $n^{f_{0}(D^{\prime})-\Sigma_{i=1}^{\infty}f_{i}(D^{\prime})\alpha_{i}}=o(n^{f_{0}(D)-\Sigma_{i=1}^{\infty}f_{i}(D)\alpha_{i}})$ as $\alpha_{l}>0$. ∎
3 Cup Product for $k=1$
Proof of theorem Theorem 1.8 for $k=1$.
2-simplicies can be added to a strongly connected 2-component in four ways (illustrated in Example 2.15):
  1. 1.
    Adding a vertex and connecting it to an existing egde.
  2. 2.
    Connecting a vertex and an unrelated edge in $A_{i}$.
  3. 3.
    A ”horn-filling”- adding an edge and a triangle to 2 existing adjacent edges.
  4. 4.
    Adding a triangle along an existing boundary of one.
Lemma 3.1.
In the range where $S^{1}_{1}(\{\alpha_{i}\}_{i})=2\alpha_{1}+\alpha_{2}>1$ and $S^{2}_{2}(\{\alpha_{i}\}_{i})=1.5\alpha_{1}+\alpha_{2}<1$, only a finite number of isomorphism types of strongly connected components can appear.
Proof.
The size of a strongly connected 2-component is bounded since all operations above decrease the expectancy of an object appearing, and the only operation of the 4 above that adds vertexes is operation (a). Assume a strongly connected 2-isomorphism type $A$ has $l+3$ vertexes. Then the expected number of the appearances of $A$ as a subcomplex of $\underline{X}(n;\alpha_{1},\alpha_{2})$ is bounded by $n^{3-3\alpha_{1}-\alpha_{2}}n^{l(1-2\alpha_{1}-\alpha_{2})}$. The first term is bounded by $n^{2}$, and the second by $n^{-\epsilon}$, where $10<\epsilon<2\alpha_{1}+\alpha_{2}-$ 1 is some positive number. Thus, the expected amount of appearances of any strongly connected 2-complex $A$ with $42/\epsilon+$ 4 vertices (of which there are finitely many options) tends to 0 as $n\to\infty$. Any strongly connected 2-complex with more than $42/\epsilon+$ 4 vertices will contain one with exactly $42/\epsilon+$ 4, and thus a.a.s would not appear as well. ∎
We now only have to address a finite collection of possible objects. $P(non-trivial\ \cup)\leq\Sigma_{C}P(C)\leq\Sigma_{C}E(C)$, where we sum over all complexes with non-trivial cup product and less than $42/\epsilon+$ 4 vertexes. Since the amount of summands doesn’t depend on $n$, it is enough to show that each one of them tends to 0.
Operations (a) and (c) induce homotopy equivalences. Operation (d) is akin to gluing a 2-cell, and thus can not increase the number of generators of $\pi_{1}$. Operation (b) is akin to attaching a 1-cell, named $e_{1}$ as a path. This can only increase the minimal number of generators of $\pi_{1}$ by 1, because of van Kampen: cover the new complex by everything but the middle of the new edge, and a circle containing the edge.
Assume a strongly connected 2-complex $A$ requires at least $\beta$ of operation (b) to construct. Then the expected number of times $A$ appears in $X$ is no greater than $n^{2-\beta(2\alpha_{1}+\alpha_{2})}=o(n^{2-\beta})$ since $12\alpha_{1}+\alpha_{2}>$ 1 in our range. Thus, only complexes where $\beta=0,1$ can appear in our range. This means that $\pi_{1}(A)$, and thus $H_{1}(A)$, have at most 1 generator. By the universal coefficient theorem, $H_{1}(A)\simeq H^{1}(A)$, since $H_{0}(A)$ is always free, and thus $H^{1}(A)$ also has at most a singe generator.
Remark 3.2.
Since for $c_{1},c_{2}\in H^{1}$ $c_{1}\cup c_{2}=(-1)^{1\cdot 1}c_{2}\cup c_{1}=-c_{2}\cup c_{1}$, this finishes the proof for rings where $2$ is not a 0-divisor.
To have $H^{1}\neq 0$ one needs a non-trivial $\pi_{1}$, so we may assume $\beta=1$. Define $v\in A$ to be an internal vertex if its link in $A$ contains a cycle. In particular, it is contained in at least 3 triangles. An internal vertex may be gotten either by applying operation (c) or (d). Operation (c) ”costs” $n^{-\alpha_{1}-\alpha_{2}}<n^{1/2}$, and thus may only be preformed once (in addition to operation (b)). Operation (d) does not have such a limitation, however the last edge of the triangle along which (d) is preformed has to be added by either operation (c) or (b), since we are adding an edge but not a vertex. If it is operation (b),
then any path from the left unmarked triangle to the right unmarked triangle is homotopic to a path outside the marked triangles. Since the complex was simply connected before adding the marked triangles, it returns to be so after their addition, and thus the strong connectivity component will have a trivial $\pi_{1}$ and thus no cup product. Otherwise, we may only have 4 internal vertices- 1 from operation (c) and 3 from the single application of operation (d) it enables. In addition there might be 1 vertex with a non-connected link (created by operation (b)).
Lemma 3.3.
Let $X$ be a 2-dimensional strongly connected complex, and assume $v$ is an external vertex of $X$. Then $X$ has a non-trivial cup product if and only if $X\setminus v$ has a non-trivial cup product.
Proof.
The link of an external vertex in a 2-dimensional complex may be either contractible or a disjoint union of contractible components. Erasing a vertex with a contractible link doesn’t effect the homotopy type of the complex. This is because any vertex is a cone above its link, and if the link is contractible then it may safely be erased along with all simplicies containing it.
Additionally, a vertex with disconnected link may be ”separated” (see Figure 1) into multiple vertices connected by edges, which is equivalent to the complex after the separation wedge with some circles (which as we saw before are not relevant to the cup-product question).
Figure 1: An example of the process of separation.
All in all, we may restrict to the subcomplex induced on the internal vertices, of which we have at most 4. This may only be (a boundary of) a tetrahedron, which plainly doesn’t have a cup product in any coefficients.
4 Cup Product for higher $k$
Proposition 4.1.
Assume we are in the range where $S_{1}^{k}(\{\alpha_{i}\}_{i})\geq 1$ and $S_{2}^{2k}(\{\alpha_{i}\}_{i})<2k+2$. Then a.a.s no strong connectivity component has more than $2k+1+a(k)$ vertexes, where $a(2)=4,a(3)=3$ and $a(k)=2,k\geq 4$.
Proof.
We have $\Sigma_{i=1}^{k+1}\binom{k+1}{i}\alpha_{i}\geq 1$ and $\Sigma_{i=1}^{2k}\binom{2k+2}{i+1}\alpha_{i}<2k+2$. A strongly-connected component of dimension $2k$ starts with a ”budget” of $n^{2k+1-\Sigma_{i=1}^{2k}\binom{2k+1}{i+1}\alpha_{i}}$, as this is the expected amount of $2k$-dim simplexes that appear. Using $\Sigma_{i=1}^{k+1}\binom{k+1}{i}\alpha_{i}\geq 1$, we see that:
$\Sigma_{i=1}^{2k}\binom{2k+1}{i+1}\alpha_{i}\geq\frac{\binom{2k+1}{2}}{k+1}(\Sigma_{i=1}^{k+1}\binom{k+1}{i}\alpha_{i})+\Sigma_{i=k+2}^{2k+1}\binom{2k+1}{i+1}\alpha_{i}>\frac{\binom{2k+1}{2}}{k+1}=\frac{(2k+1)k}{k+1}=2k+1-2+\frac{1}{k+1}$
where the first inequality is since it holds coefficient-wise ($k\geq 2$). This means that our initial budget is less than $2-\frac{1}{k+1}$. Additionally, the cost of adding a vertex is $n^{1-\Sigma_{i=1}^{2k}\binom{2k}{i}\alpha_{i}}$. By the same inequality, we know that
$\Sigma_{i=1}^{2k}\binom{2k}{i}\alpha_{i}\geq\frac{2k}{k+1}(\Sigma_{i=1}^{k+1}\binom{k+1}{i}\alpha_{i})\geq\frac{2k}{k+1}=2-\frac{2}{k+1}$
thus, adding a vertex costs at least $11-(2-\frac{2}{k+1})=\frac{2}{k+1}-$ 1. This means in turn that the amount of additional vertexes we can have is very restricted when $k\geq 2$. specifically, for $k=2$ we may have at most 4 additional vertexes, for $k=3$ only 3, and for $k\geq 4$ only 2. ∎
Proof of Theorem 1.8 for $k\geq 2$.
We have bounded the possible amounts of vertexes in strongly connected $2k$-components in proposition Proposition 4.1, and now only need to address the small amount of possible topologies that may be supported on so few vertexes. Dividing into cases by number of additional vertexes:
  1. 0,1:
    Here the only possible homotopy types are $*,S^{2k}$, both of which have cup length of 1.
  2. 2:
    If 2 vertexes were added, we see that we have a budget of less than
    $2-\frac{1}{k+1}+2(\frac{2}{k+1}-1)=\frac{3}{k+1}$
    The cost of an edge is at least $\frac{1}{k+1}$, and thus we may add at most 2 edges. Any $2k$-dim complex on $32k+$ 3 vertexes has at most $\binom{2k+3}{2}=\frac{4k^{2}+10k+6}{2}$ edges. Since we may add only 2 edges in addition to the ones added while adding vertexes, we have at most $\binom{2k+1}{2}+4k+2=\frac{4k^{2}+2k+8k+4}{2}=\frac{4k^{2}+10k+4}{2}$, and we see it is 1 less than the complete complex. This means that we have a pair of vertexes that are not connected, denote them by $v,w$.
    vw
    $lk(v)$ has $12k+$ 1 (or $2k$) vertexes, thus it is either the $(2k-1)$-dim sphere or contractible. If it’s contractible, erasing $v$ is homotopy-invariant. Similarly for $w$. If one of $v,w$ was erased- we return to the case where at most 1 vertex was added. otherwise, the only remaining question is whether the $12k+$ 1 vertexes other than $v,w$ form a simplex. If they do, the component is homotopic to a wedge of 2 spheres. Otherwise to a sphere. In both cases, there is no cup product.
  3. 3:
    In this case we have a budget of $12-\frac{1}{k+1}+3(\frac{2}{k+1}-1)=\frac{5}{k+1}-$ 1, thus this case is only relevant for $k=2,3$, for which the budget is less than $\frac{2}{3},\frac{1}{4}$ respectively. This means we may add at most 1 edge (denote it by $e$) after adding the vertexes (see Lemma 2.17). Assume another simplex is added. This operation can’t involve adding vertexes or edges, thus must be supported on $12k+$ 1 vertexes whose 1-skeleton forms a clique. This is only possible if we include $e$, otherwise we are working with an existing simplex (and if $e$ wasn’t added, the component remains contractible). Denote by $x,y$ the endpoints of $e$. Any additional operation must include $x,y$ and only their common neighbors. How many of those exist? Assume WLOG that $y$ was added after $x$. At this point, they may have at most $2k$ common neighbors (as this is the amount of neighbors of $y$). Any vertex added later may not be a common neighbor, since a simplex may be glued to a $12k-$ 1 face, which before adding $e$ can’t contain both $x,y$ (by Lemma 2.17 we may assume that the vertexes were added before the edges). Since any additional operations will only be preformed on $22k+$ 2 vertexes, the remaining 2 have contractible link, and thus can be erased without changing the homotopy type. This returns us to the case of only 1 additional vertex.
  4. 4:
    This is only relevant for $k=2$, where we remain with a budget of less than $\frac{1}{3}$. Thus no edge can be added, and thus no other simplex of any dimension. Therefore, the component is contractible as we saw before.
Remark 4.2.
To avoid confusion it is important to stress that the proposition only addresses strongly connected 2k-components. The proof that $H^{l}\neq 0$ for (in our case) $k\leq l\leq 2k$ in [Fowler19] involves proving the existence of $\partial\Delta^{l}$ as a subcomplex of $X$ with a maximal face. This maximal face is by definition ”unwitnessed” by any strongly connected component of dimension greater than $l$.
5 Steenrod Squares and Powers
Proof of Theorem 1.9.
Lemma 5.1.
Consider $\sigma$ as before. Then there exists a strongly connected pure-dimensional simplicial complex $C_{\sigma}$ with $\sigma:H^{*}(C_{\sigma};\mathbb{F}_{p})\to H^{*+deg(\sigma)}(C_{\sigma};\mathbb{F}_{p})$ a non-trivial map.
Proof.
From [AT, p.500] we know that there exists $c\in\mathbb{N}_{\geq 0}$ such that $\sigma(\iota)$ is non-trivial in $H^{*}(K(\mathbb{F}_{p},c);\mathbb{F}_{p})$, $\iota$ being a generator of $H^{c}(K(\mathbb{F}_{p},c);\mathbb{F}_{p})$ ($c$ bounding the excess of the monomials composing $\sigma$, see again [AT] for details). Further, we know from [BERGER07, Prop. 4.10.1] that $K(\mathbb{F}_{p},c)$ has a dimension-wise finite CW model. Therefore, if we truncate above dimension $c+deg(\sigma)$, we would get a finite CW-complex with $\sigma$ acting non-trivially on cohomology (as both $\iota$ and $\sigma(\iota)$ would be unaffected by the truncation). Then, by [AT, Theorem 2C.5.] (every finite CW complex is homotopy equivalent to a finite simplicial complex), and the fact that only strong connectivity components are relevant to steenrod operations, we get the result. ∎
At this point it is important to note that there is no reason to believe that $C_{\sigma}$ appears as a subcomplex of $\underline{X}(n;\alpha_{1},...)$ with any non-negligible probability. However, the next construction allows us to modify the complex s.t it does:
Construction 5.2.
Let $C$ be a finite (strongly connected) simplicial complex. Define a new complex $C^{\prime}$ as follows: add a vertex $v$ as a cone point above $C$. Then add the simplex supported on all vertexes of $C$. This results in a complex $C^{\prime}$ homotopy equivalent to $\Sigma C$ by the following map: Let $\Sigma C$ be modeled by $(C\times[0,1]/C\times\{1\})/C\times\{0\}$. Then the map that sends $v$ to $1$ and all simplexes supported on $C$ to 0 is obviously a homotopy equivalence.
This however does not preserve purity of dimension. To achieve that, assuming $dim(C)=d$, take $\Sigma^{\prime}C=sk_{d+1}(C^{\prime})$. This is seen to preserve both purity of dimension and strong connectivity. Moreover, it preserves the existence of non-zero steenrod squares hitting the maximal dimension (or lower). This is since suspension preserves them, and erasing the higher dimensional simplexes only expands the set of cocycles in relevant dimensions, thus satisfying the requirements in the corollary to Lemma 2.7.
Notice that if $C$ has complete $l$ skeleton, $\Sigma^{\prime}C$ has a complete $l+1$-skeleton. In particular, $\Sigma^{\prime}C$ always has a complete 1-skeleton if $dim(C)\geq 1$.
We now apply the construction to $C_{\sigma}$:
Lemma 5.3.
For any $C_{\sigma}$ as in Lemma 5.1 there exist $\{\alpha_{i}\}_{i}$ and $r\in\mathbb{N}$ s.t the expected amount of subcomplexes of $\underline{X}(n;\alpha_{1},...)$ isomorphic to $\Sigma^{\prime r}C_{\sigma}$ tends to infinity.
Proof.
Assume $dim(C_{\sigma})=d$, it has $v$ vertexes and minimal non-trivial cohomology in dimension  $k$. Then $\Sigma^{\prime}C_{\sigma}$ has $v+1$ vertexes and as we saw a complete 1-skeleton. This means the expected number of such subcomplexes is of the form $n^{v+1-\binom{v+1}{2}\alpha_{1}-\Sigma_{i=2}^{d+1}a_{i}\alpha_{i}}$, where $a_{i}$ are positive coefficients. This in turn means that we may add $S^{k+1}_{1}(\{\alpha_{i}\}_{i})$ at most $\frac{\binom{v+1}{2}}{k+2}$ times to the exponent while keeping the coefficients negative. Assuming $S^{k+1}_{1}(\{\alpha_{i}\}_{i})\geq 1$, the exponent will be positive (for sufficiently small $\{\alpha_{i}\}_{i\geq 2}$’s) provided that $\frac{\binom{v+1}{2}}{k+2}<v+1$, which is equivalent to $v<2k+4$.
Since each suspension increases both the minimal dimension of cohomology and the vertex amount by 1, there exists $r$ s.t $v+r<2(k+r+2)$ (equivalently $v<2k+r+2\Leftrightarrow r>v+2-2k$) and thus the expected amount of copies of $\Sigma^{\prime r}C_{\sigma}$ tends to infinity for a certain choice of $\alpha$’s, which we assume from now on. ∎
Lemma 5.4.
Denote by $M_{n}=\Sigma_{i\in\binom{[n]}{v+r}}\mathbb{1}_{i}$, where the indicator checks whether the induced ordered complex on the vertexes contains $\Sigma^{\prime r}C_{\sigma}$, and $v,r,k$ are as above. For $\alpha$’s s.t $E(M_{n})\to\infty$ and $S_{1}^{k+r}(\{\alpha_{i}\}_{i})\geq 1$, $X$ a.a.s contains more than $\frac{E(M_{n})}{2}$ copies of $\Sigma^{\prime r}C_{\sigma}$.
Proof.
We use the second moment method. By Chebyshov’s inequality
$P(|M_{n}-E(M_{n})|\geq\frac{E(M_{n})}{2})\leq 4\frac{Var(M_{n})}{E^{2}(M_{n})}$
Put in other words, we wish to show that $Var(M_{n})=o(E^{2}(M_{n}))$.
$Var(M_{n})=\Sigma_{i,j\in\binom{[n]}{v+r}}(E(\mathbb{1}_{i}\mathbb{1}_{j})-E(\mathbb{1}_{i})E(\mathbb{1}_{j}))$
We separate the sum by the size of the intersection $|i\cap j|:=m$, and treat each separately:
  • $Cov(\mathbb{1}_{i},\mathbb{1}_{j})=0$ when $|i\cap j|=0,1$, since there is no simplex of positive dimension  common to both.
  • If $12\leq m\leq d+r+$ 1, then $\Sigma_{|i\cap j|=m}E(\mathbb{1}_{i}\mathbb{1}_{j})$ (we address the subtracted term later) is equal to (up to a scalar bounded by $\binom{v+d+1}{m}^{2}$) the expected amount of subcomplexes isomorphic to complexes attained by gluing two copies of $\Sigma^{\prime r}C_{\sigma}$ along a subcomplex with $m$ vertexes.
    If $E(\mathbb{1}_{j})=\theta(n^{-c})$, then $\Sigma_{|i\cap j|=m}E(\mathbb{1}_{i}\mathbb{1}_{j})=\theta(n^{2(v+r)-2c+q_{m}-m})$, where $q_{m}\leq\Sigma_{l=1}^{m-1}\binom{m}{l+1}\alpha_{l}$. But since the $\alpha$’s are chosen s.t the complex has a positive expectance of $m-1$-dimension simplexes (their expected amount is larger than that of $\Sigma^{\prime r}C_{\sigma}$), we know that $q_{m}\leq\Sigma_{l=1}^{m-1}\binom{m}{l+1}\alpha_{l}-m<0$, and thus $n^{-2c+q_{m}-m}=o(n^{2(v+r)-2c})$. The amount of isomorphism types of such intersections doesn’t depend on $n$, and so this part of the sum is $o(n^{2(v+r)-2c})$.
  • If $d+r+2\leq m<v+r$, then $\mathbb{1}_{i}\cdot\mathbb{1}_{j}$ is an indicator of a strongly connected complex as $\Sigma^{\prime r}C_{\sigma}$ has, apart from 1 vertex, a complete $d+r$ skeleton. This complex contains $\Sigma^{\prime r}C_{\sigma}$, but has more vertexes than it. Thus, since in particular $S_{1}^{k+r}(\{\alpha_{i}\}_{i})\geq 1$, $\Sigma_{|i\cap j|=m}E(\mathbb{1}_{i}\mathbb{1}_{j})$ is (up to a constant) a sum over all isomorphism types where each summand is $o(n^{v+r-c})$, and the amount is independent of $n$.
  • In the case where the intersection is full, we have $E(M_{n})$ as the sum, which is certainly $o(E(M_{n})^{2})$.
  • The subtracted terms where the intersection is $\geq 2$ may be ignored since their contribution is $\theta(n^{2v+2r-2-2c})$ which tends to $\infty$ slower than the denominator.
All in all, we have that
$P(|M_{n}-E(M_{n})|\geq\frac{E(M_{n})}{2})\leq 4\frac{Var(M_{n})}{E^{2}(M_{n})}\xrightarrow{n\to\infty}0$
and we get the result. ∎
Lemma 5.5.
For positive $\alpha$’s s.t $E(M_{n})\to\infty$ and $S_{1}^{k+r}(\{\alpha_{i}\}_{i})>1$, $X$ a.a.s contains a component strongly connected w.r.t dimension $k+r$ isomorphic to $\Sigma^{\prime r}C_{\sigma}$
Proof.
As we saw, $X$ a.a.s contains $\theta(E(M_{n}))$ copies of $\Sigma^{\prime r}C_{\sigma}$ as subcomplexes. Let $D$ be a complex, strongly connected w.r.t dimension $k+r$, containing $\Sigma^{\prime r}C_{\sigma}$ as a proper subcomplex. Since $S^{k+r}_{1}(\{\alpha_{i}\}_{i})>1$, $E(N_{n})=o(E(M_{n}))$ by Proposition 2.19, $N_{n}$ being the expected amount of subcomplexes of $X$ isomorphic to $D$. Thus, by Markov’s inequality,
$P(N_{n}\geq E(M_{n})/2)\leq\frac{2E(N_{n})}{E(M_{n})}\xrightarrow[n\to\infty]{}0$
[…]
Thus, for any cocycle $c\in H^{k+r}(\Sigma^{\prime r}C_{\sigma})$ and any copy of $\Sigma^{\prime r}C_{\sigma}$ appearing in $X$ as a component strongly connected w.r.t dimension $k+r$, both $c$ and $\sigma c$ may be expanded to cocycles in $X$ (by extending them as 0 to the rest of the complex), and by applying Lemma 2.7 we get the result. ∎
6 Upper model
Definition 6.1.
Let $C$ be a simplicial complex that has a pair of simplexes $\sigma\subset\tau$ such that $\sigma$ is a maximal face of $\tau$, and no simplex contains $\sigma$ apart from $\tau$. $\sigma$ is then called a free face (of either $\tau$ or $C$). Erasing both $\sigma$ and $\tau$ from the complex is called an elementary collapse and induces an homotopy equivalence between the old and new complex. $C$ collapes onto dimension $d$ if there exists a series of elementary collapses starting with $C$ and resulting in a simplicial complex with only simplexes of dimension $d$ or lower. $C$ is called collapsible if it collapses onto a point.
Collapsiblility is a stronger condition than being null-homotopic. The following is a classic example:
Example 6.2.
The following is a triangulation of the (topological) Dunce Hat:
(figure is from [SANTAMARIAGALVIS2022103584]). Note the directions in which the edges of the triangle are glued. This topological space is contractible, yet there is no free face of any dimension, thus it is not collapsible.
Before proving Theorem 1.10, we need a few lemmas.
Lemma 6.3.
Let $Z$ be a strongly connected complex w.r.t dimension $l+1$. In $\overline{X}(n;\alpha_{1},...)$ with $\{\alpha_{i}\}_{i}$ such that $\beta\not\in\mathbb{Z}$, we have the following:
  1. 1.
    $E(Z\subset\overline{X})=O(n^{\lambda}),\lambda<l+1$. In other words, the budget is less than $l+1$.
  2. 2.
    Let $Z\cup\Delta^{m},m>l$ be the result of an operation expanding $Z$. Then $E(Z\cup\Delta^{m}\subset\overline{X})=O(E(Z\subset\overline{X})n^{e_{l+1}+v+l-m})$, where $v$ is the amount of new vertexes added by the operation. In other words, the cost of such an operation is $m-e_{l+1}-v-l$ (note that $e_{l+1}<0$ and that $m\geq v+l$).
Proof.
Observe that:
$E(\Delta^{k}\subset\overline{X})=\theta(n^{k+1}P(\overline{X}|_{[k+1]}=\Delta^{k}))=\theta(n^{k+1}(n^{-\alpha_{k}}+n^{1-\alpha_{k+1}}+...-O(n^{max_{i\neq j\geq k}(i+j-2(k+1)-\alpha_{i}-\alpha_{j})}))$
Recall from Definition 1.3 that $i+1-\alpha_{i}=\beta_{i}$, $max(\beta_{i})=\beta<l+1\leq k+1$, thus all powers in the second multiplicand are negative, and the correction term may be ignored. $n^{k+1}(n^{-\alpha_{k}}+n^{1-\alpha_{k+1}}+...)=\theta(n^{max_{i\geq k}\{i+1-\alpha_{i}\}})=\theta(n^{\gamma_{k}})$. But $\gamma_{k}\leq\beta<l+1$.
Denote $E(Z)=\zeta$. Suppose $Z\cup\Delta^{m}$ has $v$ vertexes more than $Z$. Then by a similar argument to the simplex:
$E(Z\cup\Delta^{m})=\theta(n^{v}(\Sigma_{i=0}n^{i-\alpha_{m+i}})\zeta)=\theta(n^{v-m-1+\gamma_{m}}\zeta)=\theta(n^{v-1+e_{m}}\zeta)$
But $e_{m}\leq l+1-m+e_{l+1}$ and we get the desired result. ∎
Part 2 of the lemma should be interpreted in the following way- the new simplex must intersect with the old complex at least at $l+1$ vertexes in order to maintain strong connectivity. If all other vertexes are new- the opperation costs $e_{l+1}$, and we pay $1$ for any additional vertex of intersection.
Lemma 6.4.
Let $C_{0}$ be a component strongly connected w.r.t dimension $k$ in a complex $C$. $C_{0}$ collapses onto dimension  $k-1$ if and only if $C_{0}$ has a growth process where no $\partial(\Delta^{d})$ for $d\geq k$ is filled.
Proof.
[…]
The other direction is obtained by retracing the steps of the collapse in the opposite order as expansion operations. ∎
Proof of 1.10.
In Lemma 6.3 we saw that an expansion operation w.r.t dimension $l+1$ costs at least $e_{l+1}$, and we start with a budget of less than $l+1$. From this we can deduce by arguments identical to those of Lemma 3.1 that only a finite number of isomorphism types of complexes strongly connected w.r.t dimension $l+1$ may appear as subcomplexes of $\overline{X}$. $P(\partial(\Delta^{d})\ filled,d\geq l+1)\leq\Sigma_{Z}P(Z)\leq\Sigma_{Z}E(Z)$, where we sum over all complexes where filling $\partial(\Delta^{d})$ is necessary in the growing possess and having less than $l+2+\frac{l+1}{e_{l+1}}$ vertexes. This means that we may prove that any complex strongly connected w.r.t dimension $l+1$ not satisfying the condition of Lemma 6.4 a.a.s doesn’t appear as a subcomplex of $\overline{X}$, and it will follow that nothing from collection a.a.s appears.
So for any choice of $k\geq l$ we need to prove that we do not fill $\partial(\Delta^{k+1})$ in components strongly connected w.r.t dimension $l+1$. Choose $k+2$ vertexes (denote $[k+2]$ for convenience), and suppose that 2 of the faces ($\hat{1}$ and $\hat{2}$, the faces missing the vertexes $1$ and $2$ respectively) are already contained in a component strongly connected w.r.t dimension  $k+1$. In order to get all of $\partial(\Delta^{k+1})$ on $[k+2]$ (without the interior) we need to fill an additional $k$ faces. Suppose we are trying to add $\widehat{k+2}$ via an expansion operation. We now divide into cases:
$k=l:$
$\widehat{l+2}$ has $l+1$ vertexes in this case, and so is itself $l$-dimensional. Adding $\widehat{l+2}$ by itself would not preserve strong connectivity w.r.t dimension $l+1$, and thus any expansion operation adding $\widehat{l+2}$ needs to be supported on $[l+1]$, and at least 1 additional vertex contained in an $l$-face already in the component. Therefore, if the operation adds a simplex of dimension $m$, it may add at most $m+1-l-2$ vertexes, and costs at least $-e_{l+1}+m-l-m-1+l+2=-e_{l+1}+1\geq 1$ by Lemma 6.3.
$k>l:$
$\widehat{k+2}$ has $k+1>l+1$ vertexes in this case, and thus an expansion of dimension $m$ adds at most $m+1-k-1\leq m+1-l-2$ vertexes, and costs at least $1$ for the same reason as in the previous case.
All in all, to obtain $\partial(\Delta^{k+1})$ as a subcomplex of the component, we need to preform at least $k\geq l$ operations with a cost of $1$ or more. Moreover, any operation filling $\partial(\Delta^{k+1})$ adds no more than $m+1-k-2\geq m+1-l-2$ new vertexes to the component, and so costs at least $1$ for the same reason as before. As we saw in Lemma 6.3 we start with a budget of less than $l+1$, and thus any component requiring filling a $\partial(\Delta^{k+1}),k\geq l$ can’t appear as a subcomplex. ∎
Remark 6.5.
The dif and only iference between the 2 cases may be illustrated with the following example:
In both diagrams, our aim is to construct the boundary of the empty triangle in the center. The left-hand diagram only has to be strongly-connected w.r.t dimension 1, and would correspond to the case of $l=0,k=1$ in the proof. The right-hand diagram, however, has to be strongly-connected w.r.t dimension 2, corresponding to the case where $l=k=1$.
In both cases $k=1$, but on the left the edges of the middle-triangle can be added ”ex nihilo” as any new simplex is required to intersect the preceding complex only along a vertex. On the right, an expansion operation is required to intersect the complex along an existing edge. This constraint is what makes these operations ”costly”.
7 Open Problems
  • In Theorem 1.9 we only address the question of whether the steenrod operations are 0 or not. We might ask a more quantitative question about the behaviour of their rank. Certainly it tends to $\infty$ (as we proved), and certainly their ”relative rank” ($\frac{rank(\sigma)}{dim(H^{d}(X))}$), for any $d$ s.t $S^{d}_{1}(\{\alpha_{i}\}_{i})\geq 1$, tends to 0. This is since, from similar arguments to the proof above, the asymptotically largest contributor to cohomology for such $d$ is always the simplest, which is empty simplexes with a maximal face.
    For the question of asymptotic rate of growth of the rank, we gave an answer depending on minimal triangulations of spaces with non-zero $\sigma$, for which the author is not aware of any mention in the literature. The answer to the rate of convergence of $\frac{rank(\sigma)}{dim(H^{d}(X))}$ to 0 should then follow from the prevalence of this minimal triangulation as opposed to that of empty simplexes with a maximal face.
  • The results in this paper do not cover all possible parameters. For the cup product, the range of $\alpha$’s discussed in Theorem 1.8 doesn’t include the case where $S^{k}_{1}(\{\alpha_{i}\})=1$, nor does it include the case where some $\alpha$’s are 0 (or approach 0 slower than $n^{-\alpha}$, for example $\frac{1}{log(n)}$).
    As for Section 5, we only proved that the algebra acts non-trivially for some range of $\alpha$’s, but have not commented about what happens in the complement. Further, we did not say the exact dimensions for which the image of $\sigma$ is non-zero (finding the minimal such dimension is an interesting question), but only proved an upper bound depending on (again) the minimal triangulation of a complex with non-zero $\sigma$.
    As for $\overline{X}$, we did not address the case where $\beta\in\mathbb{N}$. This is where the complex ”transitions” between the regime where it has complete $l-1$ skeleton to one where it has complete $l$ skeleton, thus we might see two consecutive non-trivial homologies.
  • Similar questions about other models of random simplicial complexes is left for future publications. Further, we do not address the appearance of non-zero secondary, tertiary and so on operations (such as the Massey product of various lengths).
[…]